Hw21 4.1: Problem 7 Previous Problem List Next (1 point) Get hele entering answer See a similar example (PDF) andA-Pe to solve the exercises below. Use the compound interest formulas A Round your answer to the nearest cent and make sure to include the dollar sign Find the accumulated value of an investment of $11070 for 7 years at an annual interest rate of 5.5% if the money is compounded seminannually: compounded monthly: compounded quarterly compounded continuously: Note: You can earn partial credit on this problem.
Answer #1

Solution :

Let

P- $11070

\mathrm{t = 7\: years}

\mathrm{r\:=\:0.055\:or\:5.5\%\:annual\:interest\:rate}

\mathrm{A\:=\:future\:value\:\left(accumulated\:value\right)}

\mathrm{Now,}

\mathrm{For\:compounded\:semiannually\::}

\mathrm{n = 2}

\mathrm{A=P\left(1+\frac{r}{n}\right)^{nt}}=11070\left(1+\frac{0.055}{2}\right)^{2*7}

=11070\left(1+0.0275\right)^{14}

=11070*1.0275^{14}

=11070*1.46199

=16184.23

\mathrm{and}

\mathrm{For\:compounded\:monthly\::}

\mathrm{n = 12}

\mathrm{A=P\left(1+\frac{r}{n}\right)^{nt}}

=11070\left(1+\frac{0.055}{12}\right)^{12*7}

=11070\left(1+0.00458 \right)^{84}

=11070*(1.00458)^{84}

=11070*1.46791

=16249.76

\mathrm{and}

\mathrm{For\:compounded\:quaterly\::}

\mathrm{n = 4}

\mathrm{A=P\left(1+\frac{r}{n}\right)^{nt}}

=11070\left(1+\frac{0.055}{4}\right)^{4*7}

11070 (1 0.01375)2

11070(1.01375)

=11070*1.46576

=16225.96

\mathrm{and}

\mathrm{For\:compounded\:continuously\::}

\mathrm{A=Pe^{rt}}=11070e^{0.055*7}=11070e^{0.385}=11070*1.46961=16268.58